the measure of the complement of an angle is 24 more that one fourth the measure of the supplement of the angle. find the measure of the angle.
thanks

Answer by mananth(16949)

Let the angle be x
The compliment of the angle is 90-x
The supplement of the angle is 180-x
One fourth the measure = ¼(180-x)


90-x= ¼(180-x) + 24
90-x-24 = ¼(180-x)
4(90-x-24)= 180-x
360-4x-96 = 180-x
360-96-180=4x-x
84=3x
X=28.deg
M Ananth---- mananth@hotmail.com